A motor exerts torque on a carousel for it to attain a speed of 2.5 rev/s starting from rest in 3.5 s. Find the work done by the motor if the carousel has a radius of 27 m and mass of 1.75 x10^5kg. Consider


the carousel to be a thin-walled cylinder rotating about its center.

1
Expert's answer
2021-11-21T17:32:09-0500

"\text{for thin-walled cylinder:}"

"I = mr^2"

"I = 1.75*10^5*27^2\approx1.28*10^8\ kg*m^2"

"W = I\phi,\text{where }\phi -\text{the rotation angle of the carousel in radians}"

"w_0 = 0"

"w_1= 2.5\frac{rev}{s}=5\pi\frac{rad}{s}"

"t = 3.5s"

"\epsilon= \frac{w_1-w_0}{t}=4.49\frac{rad}{s^2}"

"\phi = w_0t+\frac{\epsilon t^2}{2}= \frac{4.49*3.5^2}{2}=27.50\ rad"

"W = I\phi= 1.28*10^8*27.50 = 3.52*10^9 J"


"\text{Answer: }3.52*10^9 J"